I would really appreciate your help in the following question.
Lora has a gift card where all downloads for music cost the same amount. after downloading 2 songs, the balance on her card was $18.10. after downloading a total of 5 songs, the balance was $15.25
1. write an equation in slope-intercept form that represents the amount in dollars remaining on the card as a function of songs downloaded.
2. identify the slope of the line and tell what the slope represents.
3.identify the y intercept of the line and tell what it represents.
4. how many additional songs can Lora download when there is $15.25 left on the card?

Answer by Theo(13342)

let x = number of songs that are downloaded.

when x = 2, the remaining balance is 18.10

when x = 5, the remaining balance is 15.25

the slope intercept form of the equation for a straight line is y = mx + b

m is the slope
b is the y-intercept.

the formula for the slope is:

m = (y2 - y1) / (x2 - x1)

(x1,y1) = (2,18.10)
(x2,y2) = (5,15.25)

the x-value of the coordinate pair is the number of songs that were downloaded.
the y-value of the coordinate pair is the remaining balance.

slope = m = (y2-y1) / (x2-x1) = (15.25 - 18.10) / (5 - 2) = -2.85 / 3 = -.95

replace m with -.95 in the slope intercept form of the equation to get:

y = -.95 * x + b

now you want to solve for b, which is the y-intercept.

take any one of the point pairs on the line and replace y with the y-coordinate and replace x with the x-coordinate and solve for b.

we'll use the point (2,18.1)

y = -.95 * x + b becomes:

18.1 = -.95 * 2 + b

simplify to get:

18.1 = -1.9 + b

add 1.9 to both sides of the equation to get:

18.1 + 1.9 = -1.9 + 1.9 + b

combine like terms to get:

20 = b

that's your y-intercept.

the slope intercept form of the equation becomes:

y = -.95 * x + 20

answer to question 1:

the equation is y = -.95 * x + 20

answer to question 2:

the y-intercept is 20.

it represents the remaining balance of the account before any songs were downloaded.

answer to question 3:

find out the x-intercept.

the x-intercept is the value of x when y is equal to 0.

this will be when there is no more money remaining in the account.

find this by setting y = 0 and solving for x.

equation of y = -.95 * x + 20 becomes:

0 = -.95 * x + 20

subtract 20 from both sides of this equation to get:

-20 = -.95 * x + 20 - 20

combine like terms to get:

-20 = -.95 * x

divide both sides of this equation by -.95 to get:

-20 / -.95 = x

solve for x to get:

x = 21.0505263158 which we can rounded off to 21.05.

when she bought 21 cd, she had a small balance left.

that balance is y = -.95 * 21 + 20 = .05 dollars which is equal to 5 cents.

there's not enough money left in the account to buy another song, so this is her limit.

when she had 15.25 remaining balance, she had already bought 5 songs.

she will run out of money when she buys 21 songs.

she can buy an additional 21 - 5 = 16 songs before she runs out of money.

this can all be represented on the graph, which is shown below: